# Learning Encoding-Decoding Direction Pairs to Unveil Concepts of Influence in Deep Vision Networks

### v1

@@ 7.4.3 Evaluation of the Encoding Directions: Influence
In the context of this evaluation, we relate concept influence to the sensitivity of the network with respect to that concept when making class predictions. We assess the ability of our method to identify influential concepts to model predictions using RCAV Pfau et al. (2020). As we already mentioned in Section 3.5, RCAV originally used CAVs to accomplish this. Here we follow the proposition of Pahde et al. (2024) and use the learned encoding directions in their place. For sensitivity scores, we spatially replicate the learned encoding directions across all spatial locations in the image representation. Direction significance is tested with RCAV’s label permutation test to generate random directions, with the significance threshold set to $0.05$ and Bonferroni correction. For RCAV’s hyper-parameter $\alpha$ we use $\alpha=5.0$, as originally suggested. In the ablation studies, two metrics summarize the results: Significant Direction Count (SDC) and Significant Class-Direction Pairs (SCDP). SDC represents the number of signal vectors that significantly influence at least one model class, while SCDP tallies class-direction pairs where the signal vectors significantly affect the class. In those metrics significance is measured by taking into account the influence of random directions on the network predictions as it was already discussed in Section 3.5. In studies without $\mathcal{L}^{cur}$, we report influence metrics for signal vectors that were estimated post learning the directions with the subsample strategy discussed in Section 4.3. Let also $S_{C,i,k}$ denote the sensitivity of the model to concept $i$ when predicting class $k$ (as measured by RCAV and scaled in the range [1.0,1.0][-1.0,1.0]). When comparing against the unsupervised state-of-the-art, we use the following influence metric $\mathcal{I}^{1}$ which is defined as:
@@
i.e. the average sensitivity of the model across all concepts and classes.
7.5 Interpretability and Influence Comparison with the Unsupervised State-of-the-Art
For comparing with the unsupervised state-of-the-art in direction learning, we consider baselines based on PCA and NMF which have been previously used in Zhang et al. (2021); Fel et al. (2023c) and Graziani et al. (2023a; b). While these previous works have demonstrated the effectiveness of PCA and NMF to reveal interpretable directions in the latent space of deep networks, their evaluation mostly relied on subjective human experiments. Going beyond subjective evaluation, in this work, we take a best effort approach to compare them with our method in quantitative terms of interpretability and influence. In the Appendix (Section A.11), we also compare with the methods we extend (Doumanoglou et al., 2023; 2024) but under a different protocol for a fair comparison. For each network, we consider three sizes for the clustering, namely $I=3/4D$, $I=7/8D$, and $I=D$. When $I=D$, NMF trivially resolves to the natural latent space basis, i.e. the neuron directions. When studying EfficientNet we disregard NMF, since this network uses the SiLU activation function (Ramachandran et al., 2017) which allows negative feature activations, rendering NMF non-applicable. Tables 3,4, 5, 6, 7 summarize the comparative analysis for ResNet18, EfficientNet, Inception-v3, VGG16 and ResNet50, respectively. In the following subsections, we make a detailed discussion on those tables.
@@ 7.5.1 Analysis for ResNet18 trained on Places365
Table 3: Comparative Analysis of our Encoding-Decoding Direction Pairs (EDDP) against the Unsupervised State-of-the-Art. This table summarizes metrics for ResNet18 trained on Places365. EDDP-U stands for using Unconstrained Feature Manipulation and EDDP-C for using Constrained Feature Manipulation.
@@
Similar conclusions can be drawn from the segmentation metrics, in which the concept detectors of EDDP score substantially higher than the previous approaches. For example, when $I=448$, EDDP-C achieves $46.3$ score points in $\mathcal{S}^{1}$, surpassing NMF’s $31.77$, and reaches $0.11$ mIoU points in contrast to NMF’s $0.07$. Finally, the EDDP variants also achieve the most diverse clustering, covering a variety of visual concepts, as justified by the substantially higher values of $\mathcal{S}^{2}$ compared to the rest of the approaches. Furthermore, this diversity is consistently improved with the increase in cluster count, a phenomenon that is less prominent for the rest of the state-of-the-art baselines.
@@ 7.5.2 Analysis for EfficientNet trained on ImageNet
Table 4: Comparative Analysis of our Encoding-Decoding Direction Pairs (EDDP) against the Unsupervised State-of-the-Art. This table summarizes metrics for EfficientNet (b0) trained on ImageNet. EDDP-U stands for using Unconstrained Feature Manipulation and EDDP-C for using Constrained Feature Manipulation.
@@ 7.5.3 Analysis for Inception-v3 trained on ImageNet
Table 5: Comparative Analysis of our Encoding-Decoding Direction Pairs (EDDP) against the Unsupervised State-of-the-Art. This table summarizes metrics for Inception-v3 trained on ImageNet. EDDP-U stands for using Unconstrained Feature Manipulation and EDDP-C for using Constrained Feature Manipulation.
@@ 7.5.4 Analysis for VGG16 trained on ImageNet
Table 6: Comparative Analysis of our Encoding-Decoding Direction Pairs (EDDP) against the Unsupervised State-of-the-Art. This table summarizes metrics for VGG16 trained on ImageNet. EDDP-U stands for using Unconstrained Feature Manipulation and EDDP-C for using Constrained Feature Manipulation.
@@ 7.5.5 Analysis for ResNet50 trained on Moments in Time
Table 7: Comparative Analysis of our Encoding-Decoding Direction Pairs (EDDP) against the Unsupervised State-of-the-Art. This table summarizes metrics for ResNet50 trained on Moments in Time (MiT). EDDP-U stands for using Unconstrained Feature Manipulation and EDDP-C for using Constrained Feature Manipulation.
@@ 7.5.6 Discussion
From the previous experiments, we can draw the following, mostly straightforward, conclusions: a) EDDP’s directional clusters are monosemantic, often surpassing the previous state-of-the-art by a significant margin in monosemanticity metrics. b) EDDP also excels in terms of F1 score, where EDDP-U is consistently ranked first among the competition, and EDDP-C is often ranked second. c) In terms of segmentation and specifically $\mathcal{S}^{1}$, EDDP-U is consistently ranked first, except for EfficientNet in which the clustering of PCA scores better, yet with substantially inferior performance in most of the other interpretability metrics. In the same metric, EDDP-C is often ranked second. d) In terms of discovering a diverse set of concepts ($\mathcal{S}^{2}$) the EDDP variants score the best, often by a large margin, except for one case in VGG16 where they score on par with the best method. e) In terms of influence, in all cases except VGG16, the EDDP variants are ranked last. However, in many cases, the respective score is not far behind the competition.
In most cases, and in a broader sense, we observe that interpretability is somewhat negatively correlated with influence. Less interpretable directions may represent complex interactions (i.e. correspond to less pure human-understandable concepts) and, as a result, they are more impactful on the network predictions, in the sense that they affect many output classes. On the contrary, more interpretable directions may have a more targeted impact on specific output classes, and thus score lower in the respective metric.
@@ 7.6.1 Interpretability Losses
As already discussed previously, the motivation behind the loss terms that we introduced in Section 4.2 is to sustain or improve the interpretability of the clustering when moving from an orthogonal decoding direction set to a non-orthogonal one. In Table 8 we present interpretability metrics for the cases when we learn the directions with $\mathcal{L}^{fs}$, both $\mathcal{L}^{fs}$ and $\mathcal{L}^{eac}$, or without any of them. We also provide metrics for the case of orthogonal decoding directions (marked as Ortho). The latter is equivalent to learning with the method of Doumanoglou et al. (2023) using the Augmented Lagrangian loss of Section 4.5. We provide results regarding the directions learned for ResNet18 and EfficientNet. In these experiments, we exclude any form of Uncertainty Region Alignment.
@@ 7.6.2 Uncertainty Region Alignment
In this section we study how Uncertainty Region Alignment may affect the interpretability of the clustering and the discovery of concepts with significant influence on the network predictions. For this study we consider three network architectures: ResNet18, EfficientNet, and ResNet50 trained on the datasets already discussed in Section 7.1.
@@
In Table 10 we focus on comparing between the two Uncertainty Region Alignment variants by additionally considering metrics of concept influence. For the latter, we use RCAV’s statistical significance test in which we ground our SDC and SCDP metrics (Section 7.1). Some key takeaways drawn from these experiments for using $\mathcal{L}^{cur}$ are: a) Coverage may be improved by its use, but Redundancy can be a bit worsened, b) the concept detectors’ interpretability in terms of classification metrics is slightly inferior but somewhat comparable to when using $\mathcal{L}^{uur}$, c) in two of the three cases (ResNet18 and EfficientNet) the segmentation metrics across Uncertainty Region Alignment variants are mostly comparable, but in the third case (ResNet50) $\mathcal{L}^{cur}$ leads to a more diverse clustering (improved $\mathcal{S}^{2}$) but at the cost of a lower $\mathcal{S}^{1}$ score. d) In all cases, when comparing the two uncertainty region variants, the influence metrics favor $\mathcal{L}^{cur}$. In two of the three cases (EfficientNet and ResNet50) this improvement is substantially higher than what can be achieved when using $\mathcal{L}^{uur}$. For a more complete picture regarding the effect of $\mathcal{L}^{fso}$, which is used in EDDP-C but not in EDDP-U, see also Section A.10.
@@ 7.7 Insights on Interpretability and Influence in comparison to Supervised Direction Learning
Table 11: Comparison of EDDP (with CFM) with supervised direction learning Zhou et al. (2018). Comparing between: a) standard EDDP, b) combined concept detectors (Linear-OR) c) EDDP with the thresholds of concept detectors learned with supervision in a post initial direction learning step (/w sup thres), and d) IBD: a set of classifiers learned in a supervised way. The network here is ResNet18 and $I=512$.
@@
In this section, we compare our method with supervised direction learning, aiming to draw insights regarding interpretability and influence of concept directions. For this reason, we experiment with the last convolutional layer of ResNet18 trained on Places365. We consider our method with the constrained feature manipulation strategy, and two supervised learning approaches: a) IBD (Zhou et al., 2018) for supervised learning of decoding directions (i.e. concept detectors) and b) Pattern-CAVs (PCAVs) (Pahde et al., 2024) for supervised estimation of concept encoding directions. The supervised concept detectors and PCAVs are learned for the labels assigned to EDDP’s concept detectors at the direction labeling phase (Section 3.4).
@@ 7.8 Application: Global Model Explanations via Concept Sensitivity Testing
Figures 15,40,41,42 depict concrete examples of how each concept’s signal direction impacts the ResNet18’s class predictions. Figures 16, 43 and 44 depict similar examples for ResNet50. Concepts appearing more than once correspond to different directions that have been attributed the same label by Network Dissection. Seemingly irrelevant concepts with positive influence may have three possible explanations: a) the network has some sensitivity to those concepts (as ResNet18’s top1 accuracy is 56.51% and ResNet50’s 28.4%) b) their impact might be low, since RCAV only considers the sign of the class prediction difference before and after the perturbation, regardless of its magnitude, (thus those concepts may influence the prediction class positively, but by only a small amount) and c) their label may be misleading as the respective concept detectors do not reliably predict the concept (i.e. they exhibit a low IoU score).
@@ 7.9 Application: Local Model Explanations via Concept Contribution Maps
In this section, we apply the concept contribution analysis of Section 5.2 to provide a detailed local explanation for a prediction of ResNet18 in an image of Places365. In this example, we use direction pairs learned via EDDP-C and $I=512$. The example input image (Fig. 17 top left) is correctly predicted to belong to the baseball-field class. The output class logit for this image is ${l}_{c}=19.7$, the baseline logit is ${l}_{b}^{m}=0.0003$ and the unexplainable residual of (27) is $r=0.2816$. The top rows of Figures 17 and 18 provide segmentations which were obtained by applying our concept detectors to the patch embeddings of the input and subsequently upscaled to the original resolution. Each segmentation map highlights the region of the input that belongs to the detected concept. Above each segmentation map, we provide the index $i$ of the respective concept detector (marked as unit), the name of the concept that was attributed to the concept detector by Network Dissection, and the dataset-wide Intersection over Union score of the detector when detecting the concept in the validation split of Broden. Below each segmentation map, we provide the respective Concept Contribution Map (CCM). Each spatial element $\bm{p}$} of CCM for concept $i$ corresponds to the quantity $\phi_{{\bm{p}}}^{i}$ of (28). $\phi^{i}_{{\bm{p}}}$ contributes positively to the prediction in two cases: 1) whenever $\hat{{v}}_{{\bm{p}},i}-\hat{{v}}_{{\bm{p}},i}^{b}>0$ (i.e. when the concept content in the sample is more than the concept content in the baseline point) and the Concept-Class Relation Coefficient (CCRC) ${\bm{w}}_{c}^{T}\hat{{\bm{s}}}_{i}$ is positive and 2) whenever $\hat{{v}}_{{\bm{p}},i}-\hat{{v}}_{{\bm{p}},i}^{b}<0$, (i.e. when the concept content in the sample is less than the one in the baseline) and CCRC is negative. CCMs incorporate CCRC and thus they do not necessarily highlight the regions with the concept, as the sign of CCRC is integrated into the heatmaps (for instance: Unit 390). Above each CCM we provide the name of the concept together with the contribution of the CCM to the prediction logit after additionally integrating the correction factor of (27).
@@
Figures 19 and 20 depict image-level concept contributions by breaking down the explanation logit in parts of (27). In these Figures, CCRC indicates the Concept-Class Relation Coefficient and the total equals sample concept - baseline-concept + correction, exactly as in (27), excluding the small residual term.
The concepts that affect the prediction positively the most (Fig. 20) are batters-box-s and stadium-baseball-s as well as the absence of train and forest-road-s. As we can see in the Figures, in most cases, e.g. for the concepts grass-1, batters-box-s, stadium-baseball-s, meshed, martial-arts-gym-s, cavern-indoor-s, tree, forest-road-s,ice-skating-rink-indoor-s the sign of the difference in concept contribution between the sample and the baseline does not change with the consideration of the correction factor. However, there are some cases in which the imperfect state of convergence when learning the directions has an impact on the quality of the explanation; e.g. for the concepts grass-2,sky,wall,pool-table,train,highway-s,volleyball-court-outdoor-s,airplane. In those cases, the consideration of the correction factor changes the sign of the total concept contribution. Many of the cases that fall into this last category exhibit marginal differences in concept contribution scores between the sample and the baseline; for instance: grass-2, sky, wall, train, airplane, which could justify, at least partially, this behavior.

### v2

@@ 7.4.3 Evaluation of the Encoding Directions: Influence
We use RCAV RCAV to assess the ability of our method to identify concepts influential to model predictions and relate concept influence to the network’s concept sensitivity. In the ablation studies, two metrics summarize the results: Significant Direction Count (SDC) and Significant Class-Direction Pairs (SCDP). SDC represents the number of signal vectors that significantly influence at least one model class, while SCDP tallies class-direction pairs where the signal vectors significantly affect the class. Let $S_{C,i,k}$ denote the sensitivity of the model to concept $i$ when predicting class $k$. When comparing against unsupervised baselines, we use the following influence metric $\mathcal{I}^{1}$ which is defined as:
@@
i.e., the average sensitivity of the model across all concepts and classes.
7.5 Interpretability and Influence Comparison with Established Unsupervised Baselines
The previous unsupervised state-of-the-art (PCA; SVD; NMFCAV; CRAFT) has demonstrated effectiveness in revealing interpretable directions in the latent space of deep networks. However, as outlined in Section A.12, these approaches employed a class-conditioned protocol, tying their analyses to specific class labels. In this work, we instead operate under a class-agnostic protocol, which seeks concepts across the entire latent space, independent of class. This fundamental difference in the approaches renders a direct performance comparison with previous implementations inapplicable. Therefore, we utilize PCA and NMF, the core matrix decomposition methods upon which these prior works were built, as our principal baselines for a fair, classical comparison. We also move beyond subjective human ratings typically used for evaluation, and instead develop a rigorous, best-effort quantitative framework to compare interpretability and influence with these baselines. In Section A.15, we also compare with the methods we extend (UIBE; CBE), but under a different protocol for a fair comparison. For each network, we consider three sizes for the clustering, namely $I=3/4D$, $I=7/8D$, and $I=D$. When $I=D$, NMF trivially resolves to the natural latent space basis, i.e., the neuron directions. When studying EfficientNet, we disregard NMF since this network uses the SiLU activation function (silu), which allows negative feature activations, rendering NMF non-applicable. Tables 4,5, 6, 7, and 8 summarize the comparative analysis for ResNet18, EfficientNet, Inception-v3, VGG16, and ResNet50, respectively. In the following subsections, we have a detailed discussion of those tables.
@@ 7.5.1 Analysis for ResNet18 trained on Places365
Table 4: Comparative Analysis of our Encoding-Decoding Direction Pairs (EDDP) against unsupervised baselines. This table summarizes metrics for ResNet18 trained on Places365. EDDP-U stands for using Unconstrained Feature Manipulation and EDDP-C for using Constrained Feature Manipulation.
@@
Similar conclusions can be drawn from the segmentation metrics, in which the concept detectors of EDDP score substantially higher than the previous approaches. For example, when $I=448$, EDDP-C achieves $46.3$ score points in $\mathcal{S}^{1}$, surpassing NMF’s $31.77$, and reaches $0.11$ mIoU points in contrast to NMF’s $0.07$. Finally, the EDDP variants also achieve the most diverse clustering, covering a variety of visual concepts, as justified by the substantially higher values of $\mathcal{S}^{2}$ compared to the rest of the approaches. Furthermore, this diversity is consistently improved with the increase in cluster count, a phenomenon that is less prominent for the rest of the baselines.
@@ 7.5.2 Analysis for EfficientNet trained on ImageNet
Table 5: Comparative Analysis of our Encoding-Decoding Direction Pairs (EDDP) against unsupervised baselines. This table summarizes metrics for EfficientNet (b0) trained on ImageNet. EDDP-U stands for using Unconstrained Feature Manipulation and EDDP-C for using Constrained Feature Manipulation.
@@ 7.5.3 Analysis for Inception-v3 trained on ImageNet
Table 6: Comparative Analysis of our Encoding-Decoding Direction Pairs (EDDP) against unsupervised baselines. This table summarizes metrics for Inception-v3 trained on ImageNet. EDDP-U stands for using Unconstrained Feature Manipulation and EDDP-C for using Constrained Feature Manipulation.
@@ 7.5.4 Analysis for VGG16 trained on ImageNet
Table 7: Comparative Analysis of our Encoding-Decoding Direction Pairs (EDDP) against unsupervised baselines. This table summarizes metrics for VGG16 trained on ImageNet. EDDP-U stands for using Unconstrained Feature Manipulation and EDDP-C for using Constrained Feature Manipulation.
@@ 7.5.5 Analysis for ResNet50 trained on Moments in Time
Table 8: Comparative Analysis of our Encoding-Decoding Direction Pairs (EDDP) against unsupervised baselines. This table summarizes metrics for ResNet50 trained on Moments in Time (MiT). EDDP-U stands for using Unconstrained Feature Manipulation and EDDP-C for using Constrained Feature Manipulation.
@@ 7.5.6 Discussion
From the previous experiments, we can draw the following, mostly straightforward, conclusions: a) EDDP’s directional clusters are monosemantic, often surpassing the unsupervised baselines by a significant margin. b) EDDP also excels in terms of F1 score, where EDDP-U is consistently ranked first among the competition, and EDDP-C is often ranked second. c) In terms of segmentation and specifically $\mathcal{S}^{1}$, EDDP-U is consistently ranked first, except for EfficientNet in which the clustering of PCA scores better, yet with substantially inferior performance in most of the other interpretability metrics. In the same metric, EDDP-C is often ranked second. d) In terms of discovering a diverse set of concepts ($\mathcal{S}^{2}$) the EDDP variants score the best, often by a large margin, except for one case in VGG16 where they score on par with the best method. e) In terms of influence, in all cases except VGG16, the EDDP variants are ranked last. However, in many cases, the respective score is not far behind the competition.
In most cases, and in a broader sense, we observe that interpretability is somewhat negatively correlated with influence. Less interpretable directions may represent complex interactions (i.e. correspond to less pure human-understandable concepts) and, as a result, they are more impactful on the network predictions, in the sense that they affect many output classes. On the contrary, more interpretable directions may have a more targeted impact on specific output classes (sae-scaling), and thus score lower in the respective metric.
@@ 7.6.1 Interpretability Losses
The motivation behind the loss terms that we introduced in Section 4.2 is to sustain or improve the interpretability of the clustering when moving from an orthogonal decoding direction set to a non-orthogonal one. In Table 9 we present interpretability metrics for the cases when we learn the directions with $\mathcal{L}^{fs}$, both $\mathcal{L}^{fs}$ and $\mathcal{L}^{eac}$, or without any of them. We also provide metrics for the case of orthogonal decoding directions (marked as Ortho). The latter is equivalent to learning with the method of UIBE using the Augmented Lagrangian loss of Section 4.5. We provide results regarding the directions learned for ResNet18 and EfficientNet. In these experiments, we exclude any form of Uncertainty Region Alignment.
@@ 7.6.2 Uncertainty Region Alignment
In this section, we study how Uncertainty Region Alignment may affect the interpretability of clustering and the discovery of concepts with significant influence. For this study, we consider three network architectures: ResNet18, EfficientNet, and ResNet50.
@@
In Table 11 we focus on comparing between the two Uncertainty Region Alignment variants by additionally considering metrics of concept influence. Some key takeaways drawn from these experiments for using $\mathcal{L}^{cur}$ are: a) Coverage may be improved by its use, but Redundancy can be a bit worsened, b) the concept detectors’ interpretability in terms of classification metrics is slightly inferior but somewhat comparable to when using $\mathcal{L}^{uur}$, c) in two of the three cases (ResNet18 and EfficientNet) the segmentation metrics across Uncertainty Region Alignment variants are mostly comparable, but in the third case (ResNet50) $\mathcal{L}^{cur}$ leads to a more diverse clustering (improved $\mathcal{S}^{2}$) but at the cost of a lower $\mathcal{S}^{1}$ score. d) In all cases, when comparing the two uncertainty region variants, the influence metrics favor $\mathcal{L}^{cur}$. In two of the three cases (EfficientNet and ResNet50) this improvement is substantially higher than what can be achieved when using $\mathcal{L}^{uur}$. For a more complete picture regarding the effect of $\mathcal{L}^{fso}$, which is used in EDDP-C but not in EDDP-U, see also Section A.14.
@@ 7.7 Insights on Interpretability and Influence in comparison to Supervised Direction Learning
Table 12: Comparison of EDDP (with CFM) with supervised direction learning IBD. Comparing between: a) standard EDDP, b) combined concept detectors (Linear-OR) c) EDDP with the thresholds of concept detectors learned with supervision in a post initial direction learning step (/w sup thres), and d) IBD: a set of classifiers learned in a supervised way. The network here is ResNet18 and $I=512$.
@@
In this section, we compare our method with supervised direction learning, aiming to draw insights regarding interpretability and influence of concept directions. For this reason, we experiment with the last convolutional layer of ResNet18 trained on Places365. We consider our method with the constrained feature manipulation strategy, and two supervised learning approaches: a) IBD (IBD) for supervised learning of decoding directions (i.e. concept detectors) and b) Pattern-CAVs (PCAVs) (PCAV) for supervised estimation of concept encoding directions. The supervised concept detectors and PCAVs are learned for the labels assigned to EDDP’s concept detectors at the direction labeling phase (Section 3.4).
@@ 7.8 Application: Global Model Explanations via Concept Sensitivity Testing
Figures 15,40,41,42 depict concrete examples of how each concept’s signal direction impacts the ResNet18’s class predictions. Figures 16, 43 and 44 depict similar examples for ResNet50. Concepts appearing more than once correspond to different directions that have been attributed the same label. Seemingly irrelevant concepts with positive influence may have three possible explanations: a) the network has some sensitivity to those concepts (as ResNet18’s top1 accuracy is 56.51% and ResNet50’s 28.4%) b) their impact might be low, since RCAV only considers the sign of the class prediction difference before and after the perturbation, regardless of its magnitude, and c) their label may be misleading as the respective concept detectors do not reliably predict the concept, possibly indicated by a low IoU score.
@@ 7.9 Application: Local Model Explanations via Concept Contribution Maps
In this section, we apply the concept contribution analysis of Section 5.2 to provide a detailed local explanation for a prediction of ResNet18 in an image of Places365. In this example, we use direction pairs learned via EDDP-C and $I=512$. The example input image (Fig. 17 top left) is correctly predicted to belong to the baseball-field class. The output class logit for this image is ${l}_{c}=19.7$, the baseline logit is ${l}_{b}^{m}=0.0003$ and the unexplainable residual is $r=0.2816$. The top rows of Figures 17 and 18 provide concept segmentations which were obtained by applying our concept detectors to the patch embeddings of the input and subsequently upscaled to the original resolution. Above each segmentation map, we provide the index $i$ of the respective concept detector (marked as unit), the name of the concept attributed to the concept detector, and the dataset-wide Intersection over Union score of the detector when detecting the concept in the validation split of Broden. Below each segmentation map, we provide the respective CCM. Each spatial element of CCM for concept $i$ contributes positively to the prediction in two cases: 1) whenever the concept content in the sample is greater than the concept content in the baseline point and the CCRC is positive, and 2) whenever the concept content in the sample is less than that in the baseline and the CCRC is negative. CCMs do not necessarily highlight the regions with the concept, as the sign of CCRC is integrated into the heatmaps (for instance: Unit 390). Above each CCM we provide the name of the concept together with the contribution of the CCM to the explanation logit after additionally integrating the correction factor.
@@
Figures 19 and 20 depict image-level concept contributions by breaking down the explanation logit in components. In these Figures, the total equals sample concept - baseline-concept + correction, as discussed in Section 5.2, excluding the small residual term.
The concepts that affect the prediction positively the most (Fig. 20) are batters-box-s and stadium-baseball-s, as well as the absence of train and forest-road-s. As we can see in the Figures, in most cases, e.g. for the concepts grass-1, batters-box-s, stadium-baseball-s, meshed, martial-arts-gym-s, cavern-indoor-s, tree, forest-road-s,ice-skating-rink-indoor-s the sign of the difference in concept contribution between the sample and the baseline does not change with the consideration of the correction factor. However, there are some cases in which the imperfect state of convergence when learning the directions impacts the quality of the explanation; e.g. for the concepts grass-2,sky,wall,pool-table,train,highway-s,volleyball-court-outdoor-s,airplane. In those cases, the consideration of the correction factor changes the sign of the total concept contribution. Many of the cases that fall into this last category exhibit marginal differences in concept contribution scores between the sample and the baseline; for instance: grass-2, sky, wall, train, airplane, which could justify, at least partially, this behavior.